OpenStudy (anonymous):

What are two numbers that multiply to get -84 and add to get 17?

OpenStudy (phi):

one of them is x and the other is y

OpenStudy (phi):

xy= -84 x+y=17

OpenStudy (phi):

y= -84/x (comes from the first equation) use this in the 2nd equation: x + -84/x =17 can you solve this?

OpenStudy (anonymous):

Thank Youzzzz!